London DC Issues (Resolved)
  • Priority - Critical
  • Affecting Other - UK Servers
  • We are aware of issues affecting our London datacenter. We're currently investigating this and will update this ticket as soon as we have more information.

    Update: The issue is a networking problem caused by a set of switches. Networking engineers are currently working on the issue and servers are expected to return within the next 15 minutes.

    18:45: All servers are now back up and running. A full RFO will be posted once we have more information from the datacenter.

  • Date - 13/08/2015 18:06 - 13/08/2015 19:28
  • Last Updated - 13/08/2015 19:28
Selene Offline (Resolved)
  • Priority - Critical
  • Affecting Server - selene
  • Selene is currently offline, the cause of this is being investigated and more details will be provided shortly. Please bear with us.

    Update: This appears to be a compatibility issue with R1Soft and the CloudLinux kernel. We've booted to a vanilla CentOS kernel, will upgrade the CloudLinux version and perform another reboot to apply the changes.

    Update: The server is currently performing a filesystem check before booting.

    Update: We're still having issues booting this server and have diagnosed it as a hardware issue. We're still working on the problem and your patience is appreciated.

    Update: We're now back online but the server may perform slower than normal while a MySQL check and repair is performed.

  • Date - 23/10/2013 16:30 - 25/10/2013 13:01
  • Last Updated - 25/10/2013 13:01
SRV13 Migration (Resolved)
  • Priority - Low
  • Affecting Server - srv13
  • The migration for this server was due to be performed last Friday however the transfer was delayed. The migration to new hardware is currently running  and updates will be posted here as it progresses.

    Please remember no nameserver changes are required if your'e using our nameservers or custom nameservers using our IPs.

    Update: The transfer is currently 30% complete.

    Update: We're now at 50% complete.

    Update: We're now 75% complete

    Update: The bulk of the transfers are complete. We're now moving accounts with dedicated IPs and those missed by the first transfer.

    Update: All accounts have now been moved. Litespeed will be shutdown in 6 hours before the server is permanently shutdown.

  • Date - 12/08/2013 23:58 - 23/08/2013 22:06
  • Last Updated - 23/08/2013 22:06
SRV13 Kernel Update (Resolved)
  • Priority - Low
  • Affecting Server - srv13
  • At 11pm on 4th March SRV13 will be rebooted to upgrade to a new kernel providing both security and performance improvements.


    Total downtime is expected to be around 15 minutes while the server and services restart.

    Update: The reboot has been completed and server back online.

  • Date - 04/03/2013 14:15 - 04/03/2013 23:44
  • Last Updated - 04/03/2013 23:44
Server Reboot (Resolved)
  • Priority - Critical
  • Affecting Server - srv13
  • ItsmirHosting Server is currently suffering with performance problems due to processes holding on to server RAM. The only way for us to clear this is to reboot the server.


    This will be restarted at 12:20 today and around 15 minutes downtime is expected.

    The server has been rebooted and is currently running a File System Check. The server has not been rebooted for over 400 days. We will update this post with progress of the FSCK as it completes in stages.

    Update 13:30: The check is currently 25% complete.

    Update: 14:00: The check is now 75% complete.

    Update: 14:45 The server is now booting and should be online shortly.

  • Date - 17/01/2013 12:13 - 17/01/2013 23:37
  • Last Updated - 17/01/2013 23:37
Server13 Reboot (Resolved)
  • Priority - Critical
  • Affecting Server - srv13
  • The server has been rebooted and is currently running a File System Check. The server has not been rebooted for over 400 days. We will update this post with progress of the FSCK as it completes in stages.

  • Date - 17/01/2013 12:59 - 17/01/2013 13:01
  • Last Updated - 17/01/2013 13:01
KSP Offline (Resolved)
  • Priority - Critical
  • Affecting System - UK Servers
  • We're currently experiencing a loss of service at our primary POP in KSP, we're investigating this now and will report back shortly.


    Update @ 12:34 BST on 18/09/12: Our upstream have confirmed that they are having network issues their. Despite us maintaining two diverse links back to London, both of these seem to have died at the same time (KSP > THN & KSP > THE)

    Update @ 12:40 BST on 18/09/12: Our upstream and the Telehouse facilities have dispatched engineers to investigate immediately.

    We will update this notification as soon as we hear further feedback.

  • Date - 18/09/2012 12:32 - 18/09/2012 22:12
  • Last Updated - 18/09/2012 22:12
SRV13 DDoS Attack (Resolved)
  • Priority - Critical
  • Affecting Server - srv13
  • cPanel13 is currently being targetted by a DDoS attack. We're working on mitigating the attack now.

    Update: The attack has been mitigated and sites are back online.

  • Date - 09/05/2012 20:58 - 14/05/2012 21:33
  • Last Updated - 14/05/2012 21:33
NS2 and NS4 DDoS Attack (Resolved)
  • Priority - Critical
  • Affecting Other - Affecting: UK based accounts
  • It appears the attacker currently performing a DDoS attack against one of our reseller servers is determined to keep the site offline and have launched an attack towards the nameservers ns2.thewebhostserver.com and ns4.thewebhostserver.com.

    This may cause some problems in resolving domain names hosted on UK servers however ns1 and ns3 are currently unaffected.

    We're working to mitigate the attack but don't currently have an ETA.

  • Date - 18/04/2012 23:00 - 20/04/2012 15:19
  • Last Updated - 20/04/2012 15:19
SRV13 DDoS Attack (Resolved)
  • Priority - Critical
  • Affecting Server - srv13
  • We are currently experiencing a DDoS attack aimed at SRV13. We're working on mitigating the attack and identifying the target site as quickly as we can.

    Update: 20:49 - The IP being hit has been removed from the routing and the server is back up.

    Update: 21:50 - Despite the dedicated IP being removed the shared IP is experiencing connection issues. We're currently working to resolve this as quickly as we can.

    Update: 00:27 - We're still working on mitigating this attack however it's taking longer than expected due to the huge amount of traffic being received.

    Update 03:44 - We're working closely with our network ops team to mitigate the remnants of the attack.

    Update: 07:58 - Despite having identified and blocked the attackers our upstream provider is refusing the remove the nullroute from the IP address. We're trying to persuade them to remove the block as changing the shared IP address would only add further delay.

    Update: 08:42 - As we've had no luck getting the null removed we are now changing the shared IP on this server to 46.23.64.245. If you're using an external DNS service you will need to update your information. If you're using our nameservers you will not need to make any changes but it will need to propagate.

    Update: 09:35 - The config changes have been made and the DNS is syncing across the cluster.

    Update: 10:02 - The DNS cluster is now in sync but it'll take a little while for the new IP to sync to your ISP. You can try speeding this up by flushing your DNS cache: http://docs.cpanel.net/twiki/bin/view/AllDocumentation/ClearingBrowserCache

  • Date - 18/04/2012 20:38 - 20/04/2012 15:19
  • Last Updated - 20/04/2012 15:19

Server Status

Below is a real-time overview of our servers where you can check if there's any known issues.

Server Name HTTP FTP POP3 Server Load Uptime
Taurus